# DGS-Net: Distillation-Guided Gradient Surgery for CLIP Fine-Tuning in AI-Generated Image Detection

### v1

@@ 6 Conclusion
In this paper, we review and analyze existing strategies for fine-tuning CLIP in AI-generated image detection, and identify that catastrophic forgetting in current fine-tuning paradigms undermines the transferable priors of the pre-trained model. To address these challenges, we propose a novel Distillation-Guided Gradient Surgery Network (DGS-Net). Specifically, we orthogonally project the image gradients of the training network onto the complement of the harmful directions estimated from text gradients, thereby suppressing components irrelevant to the detection. In addition, we inject the coordinate-wise negative component of a frozen CLIP’s image gradients as a lightweight alignment, thus preserving the transferable priors and geometric structure established during large-scale pre-training. Together, they strike a balance between suppressing irrelevance and preserving transferable priors, substantially enhancing the generalization performance of AI-generated image detection.
[…]
Supplementary Material
7 Compared Detectors
@@
To provide a comprehensive benchmark for comparison, we introduce 11 existing state-of-the-art detection methods, including CNN-Spot (CVPR 2020) [wang2020cnn], UnivFD (CVPR 2023) [ojha2023towards], FreqNet (AAAI 2024) [tan2024frequency], NPR (CVPR 2024) [tan2024rethinking], Ladeda (arXiv 2024) [cavia2024real], AIDE (ICLR 2025) [yan2024sanity], C2P-CLIP (AAAI 2025) [tan2025c2p], DFFreq (arXiv 2025) [yan2025generalizable], SAFE (KDD 2025) [li2025improving], Effort (ICML 2025) [yan2024orthogonal] and NS-Net (arxiv 2025) [yan2025ns], details are as follow:
@@
1) CNN-Spot (CVPR 2020) [wang2020cnn]. CNN-Spot employs convolutional neural networks to detect synthetic content by analyzing common spatial artifacts in AI-generated images. It captures hierarchical features directly from pixel data, enabling the detection of generation anomalies.
@@
2) UnivFD (CVPR 2023) [ojha2023towards]. UnivFD demonstrates that CLIP can effectively extract artifacts from images. By training a classifier on these features, it achieves strong cross-generator generalization performance.
@@
3) FreqNet (AAAI 2024) [tan2024frequency]. FreqNet isolates high-frequency components of images via an FFT-based high-pass filter and introduces a frequency-domain learning block. This block transforms intermediate feature maps using FFT, applies learnable magnitude and phase adjustments, and reconstructs them with iFFT, enabling direct optimization in the frequency domain.
@@
4) NPR (CVPR 2024) [tan2024rethinking]. NPR targets structural artifacts introduced by up-sampling layers in generative models. It transforms input images into NPR maps that capture signed intensity differences between each pixel and its neighbors, explicitly revealing local dependency patterns characteristic of synthetic up-sampling operations.
@@
5) Ladeda (arxiv 2024) [cavia2024real]. LaDeDa is a patch-level deepfake detector that partitions each input image into 9 × 9 pixel patches and processes them using a BagNet-style ResNet-50 variant with its receptive field constrained to the same 9 × 9 region. The model assigns a deepfake likelihood to each patch, and the final prediction is obtained by globally pooling the patch-level scores.
@@
6) AIDE (ICLR 2025) [yan2024sanity]. AIDE combines low-level patch statistics with high-level semantics for AI-generated image detection. It employs two expert branches: a semantic branch, which leverages CLIP-ConvNeXt embeddings to detect content inconsistencies, and a patchwise branch, which selects patches by spectral energy and applies a lightweight CNN to capture fine-grained artifacts.
@@
7) C2P-CLIP (AAAI 2025) [tan2025c2p]. C2P-CLIP concludes that CLIP achieves classification by matching similar concepts rather than discerning true and false. Based on this conclusion, they propose category common prompts to fine-tune the image encoder by manually constructing category concepts combined with contrastive learning.
@@
8) DFFreq (arxiv 2025) [yan2025generalizable]. DFFreq first utilizes a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT is designed to enrich the extraction of local forgery features from different perspectives.
@@
9) SAFE (KDD 2025) [li2025improving]. SAFE replaces conventional resizing with random cropping to better preserve high-frequency details, applies data augmentations such as Color-Jitter and RandomRotation to break correlations tied to color and layout, and introduces patch-level random masking to encourage the model to focus on localized regions where synthetic pixel correlations typically emerge.
@@
10) Effort (ICML 2025) [yan2024orthogonal]. Effort find that a naively trained detector very quickly shortcuts to the seen fake patterns, collapsing the feature space into a low-ranked structure that limits expressivity and generalization. Thus, they decompose the feature space into two orthogonal subspaces, for preserving pre-trained knowledge while learning forgery.
@@
11) NS-Net (arxiv 2025) [yan2025ns]. NS-Net uses the feature homogeneity extracted by the text encoder to replace the semantic information of the features extracted by the image encoder, and uses NULL-Space to decouple the semantic information, retaining the artifact information related to the forgery detection task.
@@
CNN-Spot [wang2020cnn] 51.3 70.8 52.2 76.9 52.2 78.0 52.4 76.8 60.6 88.7 62.0 90.1 68.0 93.8 51.6 62.7 56.3 79.7
UnivFD [ojha2023towards] 64.2 81.0 84.0 92.5 84.5 92.1 84.3 92.3 82.5 91.4 82.5 91.4 73.4 84.7 69.0 80.3 78.1 88.2
FreqNet [tan2024frequency] 68.2 78.5 89.8 97.8 90.2 97.9 88.8 97.6 97.6 99.7 98.1 99.8 97.3 99.7 70.5 93.1 87.6 95.5
NPR [tan2024rethinking] 64.6 83.0 80.9 99.3 80.2 99.0 79.9 99.3 88.0 99.9 86.9 99.9 94.6 100.0 63.6 98.9 79.8 97.4
Ladeda [cavia2024real] 79.8 87.5 98.3 99.8 98.2 99.8 98.5 99.8 98.6 99.9 98.5 99.9 98.0 99.8 83.9 98.5 94.2 98.1
AIDE [yan2024sanity] 90.1 98.5 98.3 99.9 97.9 99.8 98.1 99.8 98.4 99.9 98.3 100.0 98.5 100.0 96.7 99.5 97.0 99.7
C2P-CLIP* [tan2025c2p] 80.3 98.6 94.3 99.6 93.0 99.8 89.3 99.4 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 94.6 99.7
DFFreq [yan2025generalizable] 90.3 99.2 94.8 99.2 95.7 99.2 93.9 99.0 99.3 100.0 99.2 100.0 99.2 100.0 95.2 99.5 96.0 99.5
SAFE [li2025improving] 80.7 96.0 94.8 98.5 95.2 98.4 93.2 97.9 97.2 99.9 97.4 99.9 97.1 99.9 94.7 99.1 93.8 98.7
Effort [yan2024orthogonal] 76.4 98.3 95.7 99.8 95.7 99.8 93.7 99.6 99.9 100.0 100.0 100.0 99.8 100.0 100.0 100.0 95.2 99.7
NS-Net [yan2025ns] 92.5 99.4 98.2 99.9 99.0 99.9 98.2 99.8 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.5 99.9
@@
Table 7: Cross-Diffusion-Sources Evaluation on the Diffusion Test Set of UniversalFakeDetect [ojha2023towards].
@@
Table 8: Cross-model Average Precision (A.P.) Performance on the AIGIBench [li2025artificial] Dataset.
@@
Table 9: Cross-model Accuracy (Acc.) Performance on the GenImage [zhu2024genimage] Dataset. We further experimentally verify the beneficial and harmful components of the gradient.
8 Evaluation on UniversalFakeDetect
@@
9 AP Result on AIGIBench
Follow the evaluation metrics used in baselines [ojha2023towards, tan2024frequency], which include average precision score (A.P.) and accuracy (Acc.), we add the A.P. results on AIGIBench [li2025artificial], as shown in Table 8. Furthermore, our method achieves state-of-the-art performance among the baselines in terms of AP, underscoring the superiority of the proposed approach. This indicates that our method not only excels in overall prediction accuracy but also maintains robust performance across different decision thresholds, thereby demonstrating its effectiveness in distinguishing between classes.
10 More Confirmatory Studies
As discussed in Sec.˜3 and Sec.˜4, we define the positive half-space of text gradients as harmful directions that should be suppressed, while the negative half-space of image gradients corresponds to beneficial directions that preserve the prior knowledge of the pre-trained model. To further validate this definition and theoretical reasoning, we conducted an additional experiment in which we directly used text gradients as harmful directions and treated all image gradients as beneficial directions. The results of this experiment are shown in Table 9. These results clearly demonstrate the necessity of categorizing gradients into harmful and beneficial components and treating them separately. For instance, injecting all image gradients as beneficial yields only limited improvements over the baseline, particularly on diffusion models; likewise, removing all text gradients as harmful results in performance substantially below our approach. In contrast, our method selectively projects and removes only the positive half-space of text gradients (harmful semantics) while injecting only the negative half-space of image gradients (beneficial priors). This strategy achieves the highest mAcc of 97.6%, approaching saturation on multi-class GANs and producing the most significant gains on diffusion models, thereby substantially improving cross-generator generalization.

### v2

@@ 6 Conclusion
In this paper, we review and analyze existing strategies for fine-tuning CLIP in AI-generated image detection, and identify that catastrophic forgetting in current fine-tuning paradigms undermines the transferable priors of the pre-trained model. To address these challenges, we propose a novel Distillation-Guided Gradient Surgery Network (DGS-Net). Specifically, we orthogonally project the image gradients of the training network onto the complement of the harmful directions estimated from text gradients, thereby suppressing components irrelevant to the detection. In addition, we inject the coordinate-wise negative component of a frozen CLIP’s image gradients as a lightweight alignment, preserving the transferable priors and geometric structure established during large-scale pre-training. Together, they strike a balance between suppressing irrelevance and preserving transferable priors, substantially enhancing the generalization performance of AI-generated image detection.
Impact Statement
This paper presents work aimed at advancing the field of Security in Trustworthy Machine Learning. It investigates and proposes an AI-generated image detection approach that demonstrates potential for mitigating the malicious use of generative models and may produce a positive social impact.
Appendix A Compared Detectors
@@
To provide a comprehensive benchmark for comparison, we introduce 11 existing state-of-the-art detection methods, including CNN-Spot (CVPR 2020) (wang2020cnn), UnivFD (CVPR 2023) (ojha2023towards), FreqNet (AAAI 2024) (tan2024frequency), NPR (CVPR 2024) (tan2024rethinking), Ladeda (arXiv 2024) (cavia2024real), AIDE (ICLR 2025) (yan2024sanity), C2P-CLIP (AAAI 2025) (tan2025c2p), DFFreq (arXiv 2025) (yan2025generalizable), SAFE (KDD 2025) (li2025improving), Effort (ICML 2025) (yan2024orthogonal) and NS-Net (arxiv 2025) (yan2025ns), details are as follow:
@@
1) CNN-Spot (CVPR 2020) (wang2020cnn). CNN-Spot employs convolutional neural networks to detect synthetic content by analyzing common spatial artifacts in AI-generated images. It captures hierarchical features directly from pixel data, enabling the detection of generation anomalies.
@@
2) UnivFD (CVPR 2023) (ojha2023towards). UnivFD demonstrates that CLIP can effectively extract artifacts from images. By training a classifier on these features, it achieves strong cross-generator generalization performance.
@@
3) FreqNet (AAAI 2024) (tan2024frequency). FreqNet isolates high-frequency components of images via an FFT-based high-pass filter and introduces a frequency-domain learning block. This block transforms intermediate feature maps using FFT, applies learnable magnitude and phase adjustments, and reconstructs them with iFFT, enabling direct optimization in the frequency domain.
@@
4) NPR (CVPR 2024) (tan2024rethinking). NPR targets structural artifacts introduced by up-sampling layers in generative models. It transforms input images into NPR maps that capture signed intensity differences between each pixel and its neighbors, explicitly revealing local dependency patterns characteristic of synthetic up-sampling operations.
@@
5) Ladeda (arxiv 2024) (cavia2024real). LaDeDa is a patch-level deepfake detector that partitions each input image into 9 × 9 pixel patches and processes them using a BagNet-style ResNet-50 variant with its receptive field constrained to the same 9 × 9 region. The model assigns a deepfake likelihood to each patch, and the final prediction is obtained by globally pooling the patch-level scores.
@@
6) AIDE (ICLR 2025) (yan2024sanity). AIDE combines low-level patch statistics with high-level semantics for AI-generated image detection. It employs two expert branches: a semantic branch, which leverages CLIP-ConvNeXt embeddings to detect content inconsistencies, and a patchwise branch, which selects patches by spectral energy and applies a lightweight CNN to capture fine-grained artifacts.
@@
7) C2P-CLIP (AAAI 2025) (tan2025c2p). C2P-CLIP concludes that CLIP achieves classification by matching similar concepts rather than discerning true and false. Based on this conclusion, they propose category common prompts to fine-tune the image encoder by manually constructing category concepts combined with contrastive learning.
@@
8) DFFreq (arxiv 2025) (yan2025generalizable). DFFreq first utilizes a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT is designed to enrich the extraction of local forgery features from different perspectives.
@@
9) SAFE (KDD 2025) (li2025improving). SAFE replaces conventional resizing with random cropping to better preserve high-frequency details, applies data augmentations such as Color-Jitter and RandomRotation to break correlations tied to color and layout, and introduces patch-level random masking to encourage the model to focus on localized regions where synthetic pixel correlations typically emerge.
@@
10) Effort (ICML 2025) (yan2024orthogonal). Effort find that a naively trained detector very quickly shortcuts to the seen fake patterns, collapsing the feature space into a low-ranked structure that limits expressivity and generalization. Thus, they decompose the feature space into two orthogonal subspaces, for preserving pre-trained knowledge while learning forgery.
@@
11) NS-Net (arxiv 2025) (yan2025ns). NS-Net uses the feature homogeneity extracted by the text encoder to replace the semantic information of the features extracted by the image encoder, and uses NULL-Space to decouple the semantic information, retaining the artifact information related to the forgery detection task.
Table 7: Cross-Diffusion-Sources Evaluation on the Diffusion Test Set of UniversalFakeDetect (ojha2023towards).
@@
CNN-Spot (wang2020cnn) 51.3 70.8 52.2 76.9 52.2 78.0 52.4 76.8 60.6 88.7 62.0 90.1 68.0 93.8 51.6 62.7 56.3 79.7
UnivFD (ojha2023towards) 64.2 81.0 84.0 92.5 84.5 92.1 84.3 92.3 82.5 91.4 82.5 91.4 73.4 84.7 69.0 80.3 78.1 88.2
FreqNet (tan2024frequency) 68.2 78.5 89.8 97.8 90.2 97.9 88.8 97.6 97.6 99.7 98.1 99.8 97.3 99.7 70.5 93.1 87.6 95.5
NPR (tan2024rethinking) 64.6 83.0 80.9 99.3 80.2 99.0 79.9 99.3 88.0 99.9 86.9 99.9 94.6 100.0 63.6 98.9 79.8 97.4
Ladeda (cavia2024real) 79.8 87.5 98.3 99.8 98.2 99.8 98.5 99.8 98.6 99.9 98.5 99.9 98.0 99.8 83.9 98.5 94.2 98.1
AIDE (yan2024sanity) 90.1 98.5 98.3 99.9 97.9 99.8 98.1 99.8 98.4 99.9 98.3 100.0 98.5 100.0 96.7 99.5 97.0 99.7
C2P-CLIP* (tan2025c2p) 80.3 98.6 94.3 99.6 93.0 99.8 89.3 99.4 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 94.6 99.7
DFFreq (yan2025generalizable) 90.3 99.2 94.8 99.2 95.7 99.2 93.9 99.0 99.3 100.0 99.2 100.0 99.2 100.0 95.2 99.5 96.0 99.5
SAFE (li2025improving) 80.7 96.0 94.8 98.5 95.2 98.4 93.2 97.9 97.2 99.9 97.4 99.9 97.1 99.9 94.7 99.1 93.8 98.7
Effort (yan2024orthogonal) 76.4 98.3 95.7 99.8 95.7 99.8 93.7 99.6 99.9 100.0 100.0 100.0 99.8 100.0 100.0 100.0 95.2 99.7
NS-Net (yan2025ns) 92.5 99.4 98.2 99.9 99.0 99.9 98.2 99.8 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.5 99.9
@@
Table 8: Cross-model Average Precision (A.P.) Performance on the AIGIBench (li2025artificial) Dataset.
@@
Table 9: Cross-model Accuracy (Acc.) Performance on the AIGCDetectBench (zhong2023patchcraft) Dataset. We further experimentally verify the beneficial and harmful components of the gradient.
@@
Appendix B Evaluation on UniversalFakeDetect
@@
Appendix C AP Result on AIGIBench
Follow the evaluation metrics used in baselines (ojha2023towards; tan2024frequency), which include average precision score (A.P.) and accuracy (Acc.), we add the A.P. results on AIGIBench (li2025artificial), as shown in Table 8. Furthermore, our method achieves state-of-the-art performance among the baselines in terms of AP, underscoring the superiority of the proposed approach. This indicates that our method not only excels in overall prediction accuracy but also maintains robust performance across different decision thresholds, thereby demonstrating its effectiveness in distinguishing between classes.
Appendix D More Confirmatory Studies
As discussed in Section˜3 and Section˜4, we define the positive half-space of text gradients as harmful directions that should be suppressed, while the negative half-space of image gradients corresponds to beneficial directions that preserve the prior knowledge of the pre-trained model. To further validate this definition and theoretical reasoning, we conducted an additional experiment in which we directly used text gradients as harmful directions and treated all image gradients as beneficial directions. The results of this experiment are shown in Table 9. These results clearly demonstrate the necessity of categorizing gradients into harmful and beneficial components and treating them separately. For instance, injecting all image gradients as beneficial yields only limited improvements over the baseline, particularly on diffusion models; likewise, removing all text gradients as harmful results in performance substantially below our approach. In contrast, our method selectively projects and removes only the positive half-space of text gradients (harmful semantics) while injecting only the negative half-space of image gradients (beneficial priors). This strategy achieves the highest mAcc of 97.6%, approaching saturation on multi-class GANs and producing the most significant gains on diffusion models, thereby substantially improving cross-generator generalization.